OpenStudy (anonymous):

Help me with this assignment please.. Solving Systems of Linear Equation by Subtitution 3a+6b= -10 b=-2a-5

OpenStudy (anonymous):

replace b in equation 1 by \[-2a-5\] and don't forget the parentheses you get \[3a+6(-2a-5)=-10\]

OpenStudy (anonymous):

latex not working?

OpenStudy (angela210793):

3a+6b= -10 replace -2a-5 instead of b in the 1st eq b=-2a-5 3a+6(-2a-5)=-10 3a-12a-30+10=0 -9a=20a=-20/9 b=-2*(-20/9)-5

OpenStudy (anonymous):

substitute -2a-5 into the first equation and it will become 3a+6(-2a-5)=-10 3a-12a-30=-10 -9a=20 a=-20/9

OpenStudy (anonymous):

how about for the "b"..?

OpenStudy (anonymous):

substitute the answer for a back to the equation, b=-2a-5